Question 1

Walter & Randstad. Ltd. is a European is a single distributor of a product that grew popularity among European consumers. Below is the income statement for the three most recent months.

Walter & Randstad.

Income Statements

For the three Months Ended December 31

October

November

December

Sales in units

5000

5500

6000

Sales

$500,000

$550,000

$600,000

Cost of Goods Sold

$300,000

$320,000

$360,000

Gross Margin

$200,000

$230,000

$240,000

Selling & Administrative Expenses:

Advertising expense

$29,000

$29,000

$29,000

Salaries & commissions

$34,000

$38,000

$40,000

Insurance expenses

$8,000

$8,000

$8,000

License Fee

$60,000

$62,000

$65,000

Depreciation expenses

$20,000

$20,000

$20,000

Total Selling & Administrative Expenses

$137,000

$143,000

$147,000

Net Operating Income

$63,000

$87,000

$97,000

Required:

  1. Identify each of the company’s expenses (including Cost of goods sold) as variable, fixed or mixed cost.
  2. Using the high-low method, identify and separate mixed expenses into variable and fixed elements. State the cost formula for each mixed expenses.                                                                                                                                                         
  3. Redo the company’s income statement at the 6000 unit level of activity using the contribution format.
Answer #1

Cost of goods sold= Mixed cost

Advertising expense= Fixed cost

Salaries & commissions= Mixed cost

Insurance expense= Fixed cost

License fees= Mixed cost

Depreciation expenses= Fixed cost

  • Using High low method -

Formula=

Variable Cost per unit = Cost at Highest Activity Units− Cost at Lowest Activity Units / Highest activity units−Lowest activity units

Fixed Cost = Activity Cost - Variable cost per unit * Activity Unit

By using the above formula we can seperate mixed cost into variable and fixed cost for cost of good sold , salaries and comissions , license fees

1. Cost of good sold

Variable cost per unit = 360000-300000 / 6000-5000

= 60000 / 1000

= 60 per unit

Variable cost = 60 * 6000 = 360000

Fixed cost :

360000 = 60*6000 + fixed cost

Fixed Cost = 0

2.Salaries And Comissions

Variable cost per unit = 40000-34000 / 6000-5000

= 6000 / 1000

=6

Variable cost = 6 * 6000 = 36000

Fixed cost:

40000 = 6 * 6000 + fixed cost

40000 - 36000 = Fixed cost

4000 = Fixed Cost

3. License Fees

Variable Cost per unit = 65000 - 60000 / 6000-5000

= 5000 / 1000

= 5

variable cost = 5 * 6000 =30000

Fixed Cost:

65000 = 6000 *5 + Fixed cost

65000 - 30000 = Fixed cost

35000 = Fixed cost

Cost formula for mixed expense =

y=a + b*x

where y = total cost

a = fixed cost

b = variable cost per unit

x = number of units

above fixed cost is calculated with same formula.

Income statement at 6000 units using contribution format

Sales ( 6000 * 100 ) 600000

- Variable cost

: variable cost of goods sold (3,60,000)

: variable license fees (5 * 6000) (30000)

: variable slaries and comission (6 * 6000) (36000)

= Contribution 174000

- Fixed cost

: fixed cost of good sold 0

: fixed license fees (35000)

: fixed salaries and comission expenses (4000)

: advertising expenses (29000)

: insurance expenses (8000)

: depericiation expenses (20000)

= Profit 78000